Which of the following is an advantage of convertible bonds? Investors can convert the bonds into...

  1. Which of the following is an advantage of convertible bonds?

    Investors can convert the bonds into higher coupon rate bonds.

    Investors can choose to hold the company's bonds or convert the bonds into its stock.

    Investors are paid a penalty on the conversion of the bonds.

    Investors are redeemed for the difference between the face value and the market price on redemption of the bonds.

    Investors can claim interest for the remaining life of the bonds on the bonds' early conversion.

5 points   

QUESTION 2

  1. If the Federal Reserve loosens money supply then:

    inflation will decrease.

    interest rates will decrease.

    sale of Treasury securities will increase.

    credit supply will decrease.

    economic activity will decrease.

5 points   

QUESTION 3

  1. A bond purchased for $950 was sold for $980 after one year. The interest received during the year is $25. The bond's yield is:

    2.23%

    5.79%

    8.12%

    5.25%

    9.36%

5 points   

QUESTION 4

  1. Investors with a _____ will demand a higher rate of return.

    higher time preference for consumption

    lower exposure to economic risks

    lower access to production opportunities

    higher financial creditworthiness

    lower default premium
Answer #1

Question 1:

Convertible bonds are bond which can be converted into stocks, the investor has the advantage to hold such bond or convert in shares.

The correct answer is option b)

Question 2:

If the Federal Reserve loosens money supply then, money becomes cheap, money supply increases and interest rate decreases.

the correct answer is option b)

Question 3:

The bond yield is equal to its holding return which is:

Interest earnings + Capital gain = 25+ (980-950) = $55

Yield = Total return / Bond purchase price

=55/950

=5.79%

The correct option is b)

Question 4:

Investors with lower exposure to economic risks will demand a higher rate of return,

the correct option is b)